OpenStudy (anonymous):

PLEASE HELP! I DON'T KNOW THIS AND IT IS A BIG GRADE. The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the acceleration due to gravity, and t = time. Solve this equation for g, and use your formula to determine the acceleration due to gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 meters.

OpenStudy (anonymous):

g = 9.8 ft/sec² ANSWER

OpenStudy (anonymous):

thank you so i would just right g=9.8 feet/seconds? @lionblood

OpenStudy (anonymous):

This might help you. http://openstudy.com/study#/updates/530bb250e4b04a7f3c1db864

OpenStudy (anonymous):

that one says feet,mine is meters @lionblood

OpenStudy (anonymous):

Hit a dilemma? ^^

OpenStudy (anonymous):

yes i have no idea how to do this @Supreme_Kurt

OpenStudy (anonymous):

Just call me Kurt ^^ Now, the reason you're getting all lost and confused is because you weren't showed how to actually solve this problem, like, step-by-step. If you'll let me, I'll be able to do just that :) Ready?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Now, start with this: \[\Large d = 0.5\color{red}gt^2\] And we need to figure out the "acceleration due to gravity" or g. To do that, we need to get g ALL ALONE on one side of the equation. Any idea how to proceed with that? :)

OpenStudy (anonymous):

multiply the sides by 0.5?

OpenStudy (anonymous):

Almost right. In time, you'll learn that "almost right" is just a nice way of saying "wrong" But that's okay. Better you make mistakes here than in the actual exams or something. This is how we learn :D If I'm not mistaken, your goal is to get rid of the 0.5 on the right-side, yes? Your heart's in the right place (hehe) But multiplying both sides by 0.5 won't do the trick. That'll just turn the 0.5 on the right-side into a 0.25. If that 0.5 were a 1 instead, that'd be as if it just disappeared, since 1 multiplied to anything gets you, well, that thing. You know what I mean. That said, what do you multiply to 0.5 so that you get 1? It sure as hell isn't 0.5 ;)

OpenStudy (anonymous):

0.5x2 would be 1

OpenStudy (anonymous):

That is correct :D Now, if we multiply both sides by 2, what happens to the equation? ^^ Show me.

OpenStudy (anonymous):

i dont know :( this makes me look stupid

OpenStudy (anonymous):

Hey, just to put this there, if you're stuck or you think you've hit a dead end, tell me immediately. I'm here to teach you, not judge you ^^ And you're not stupid unless you admit it to yourself, so do yourself a favour and don't do that, okay? Now, back to our equation. \[\Large d = 0.5\color{red}gt^2\]

OpenStudy (anonymous):

We multiply both sides by 2... nice and slowly... \[\Large \color{green}2\cdot d = \color{green}2\cdot 0.5\color{red}gt^2\] With me so far?

OpenStudy (anonymous):

so far yes

OpenStudy (anonymous):

Okay. 2d. Nothing we can do about that now. \[\Large \color{blue}{2d} = \color{green}2\cdot 0.5\color{red}gt^2\] Still with me?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Now, about this part. \[\Large 2d = \color{blue}{2\cdot 0.5}gt^2\] 2 times 0.5 What does this become? You said yourself just a while ago ;)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

That's good. And you know it's pointless to write "1" times anything right, because it doesn't really do anything? ie 1 times 5 is just 5, so might as well just write 5 instead of 1x5 etc. Getting me? :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay, I'll demonstrate. :D \[\Large 2d = \color{blue}{2\cdot0.5}gt^2\] 2 times 0.5 is just 1. \[\Large 2d = \color{blue}1gt^2\] But 1 times gt^2 is jugt gt^2 so... \[\Large 2d = gt^2\] okay? ^^

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Now, we need to get rid of that t^2. It's multiplied to g, so to remove it, you...?

OpenStudy (anonymous):

multiply?

OpenStudy (anonymous):

No... okay, to undo multiplication, you divide :D To undo division, you multiply. To undo addition, you subtract. To undo subtraction, you add. This is basic stuff, so please remember it :D To undo the multiplication of t^2, you *divide* both sides by t^2. And you get...?

OpenStudy (anonymous):

this is way to hard for me, i wanna say g^2?

OpenStudy (anonymous):

No... okay, let me show you :D \[\large 2d = gt^2\] We now divide both sides by t^2 \[\Large \frac{2d}{t^2}= \frac{gt^2}{t^2}\] What happens next?

OpenStudy (anonymous):

you solve it

OpenStudy (anonymous):

Of course -_- But specifically, what step comes next? :)

OpenStudy (anonymous):

\[\Large \frac{2d}{t^2}=\frac{g\color{blue}{t^2}}{\color{red}{t^2}}\] These two cancel out, don't they?

OpenStudy (anonymous):

ohh yes

OpenStudy (anonymous):

And what does that leave you with?

OpenStudy (anonymous):

2d over t^2 = gt?

OpenStudy (anonymous):

Why would there be a t left over? :P

OpenStudy (anonymous):

ohh sorry lol, 2d over t^2 = g

OpenStudy (anonymous):

And that's it, you've solved for g, as per the instructions of the problem. \[\Large g = \frac{2d}{t^2}\]

OpenStudy (anonymous):

Now, to solve for the actual value of g, well, we need d and t. distance and time. Any idea what those are? :)

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Two numbers are staring you right in the face. I quote your problem: "...if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 meters. " The *distance* and the *time* must be there somewhere...

OpenStudy (anonymous):

490 is distance and 10 seconds is the time

OpenStudy (anonymous):

That's correct. Now just plug those in. d is distance, t is time: \[\Large g = \frac{2(490)}{(10)^2}\] and do the math :D

OpenStudy (anonymous):

g=980 over 100?

OpenStudy (anonymous):

Yup. Simplify? :)

OpenStudy (anonymous):

9 seconds?

OpenStudy (anonymous):

nope. Too simple. Keep it precise. to divide by 100, simply move the decimal point two places to the left. ;)

OpenStudy (anonymous):

what would it be, im stuck

OpenStudy (anonymous):

@Supreme_Kurt

OpenStudy (anonymous):

980.0 move the decimal point two places to the left... to divide it by 100. and you get...?

OpenStudy (anonymous):

9.8?

OpenStudy (anonymous):

Which is indeed, what Lion gave you (sort of) Except now, you know where it came from :) It's m/s^2 btw, not feet/s^2. That's it :D

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

No problem :) Practice makes perfect, so keep right on practicing, okay? ^^